A lighter-than-air spherical balloon and its load of passengers and ballast are floating stationary above the earth. Ballast is weight (of negligible volume) that can be dropped overboard to make the balloon rise. The radius of this balloon is 5.19 m. Assuming a constant value of 1.29 kg/m3 for the density of air, determine how much weight must be dropped overboard to make the balloon rise 197 m in 12.0 s.
Answer #1

When system is in equilibrium,

Fnet = Fb - M g = 0

and Fb = Volume of fluid displaced x density of fluid x g


Fb = (4 x pi x 5.19^3 / 3) (1.29)( g)

Fb = 755.4g


As Fb = M g So total mass is 755.4 kg .


d = v0 + a t^2 /2


197 = 0 + a (12^2) / 2

a = 2.74 m/s^2

upsse mass m is dropped then

Fnet = Fb - (M - m)g = (M - m)a


755.4g - (755.4 - m)g = (755.4 - m)(2.74)

m g = 2066.86 - 2.74m

m (9.81 + 2.74) = 2066.86

m = 164.7 kg Or 1615.6 N .....ANs